The Terminal Tobacco Company estimates that each case of cigars they produce can be sold for $65. Fixed costs of production are estimated at $35,000, while variable costs are estimated at $15 per case. On the basis of this information, answer the following questions:

a. What is the firm’s break-even point in units?

b. What is the firm’s break-even point in sales dollars?

c. How much profit will the company earn from the sale of 1,500 cases?

d. How many cases could Terminal have to sell to earn $180,000 profit?
Answer #1

het x denotes no, of cases Sold Revenue = 65x Told Cost = 35000 + 15x porofit = 66x - 13000 + 15x) = Sox - 35000 (9) o Bokeve